OpenStudy (istim):

A paint store sells 270 cans of paint per month at a price $32 each. A customer survey indicates that for each $1.20 decrease in price, sales will increase by 6 cans of paint. Determine the demand, or price, function.

OpenStudy (istim):

My answer is 0.2x, but the actual answer is 86-0.2x.

OpenStudy (istim):

First, I found m thru m=1.20/6=0.2 Then I gound m by 1.20=(0.2)6+b

OpenStudy (istim):

got, not gound.

OpenStudy (istim):

Don't give a medal please! I need assistance!

OpenStudy (istim):

Please! This question is relatively simple.

OpenStudy (mimi_x3):

hm, calculus?

OpenStudy (istim):

Yes. IT should be easy.

OpenStudy (mimi_x3):

looks like a minimum/maximum problem..

OpenStudy (istim):

No, even easier than that.

OpenStudy (istim):

Please, anyone, tell me how to achieve the correct answer.

OpenStudy (mimi_x3):

lol, i think there's a formula for it..which i cant remmember.. but i found this on google: http://en.allexperts.com/q/Calculus-2063/2010/7/Rate-Change-3.htm

OpenStudy (istim):

You looked up the actual question. Huh.

OpenStudy (istim):

The real answer is 86-0.2x.

OpenStudy (istim):

I got 0.2x only.

OpenStudy (mimi_x3):

The formula to solve this is \(P(x) = R(x) -C(x)\), right?

OpenStudy (istim):

That's something else.

OpenStudy (istim):

The profit function.

OpenStudy (istim):

It's the price function here.

OpenStudy (mimi_x3):

well, whats the formula to solve this? I might be able to work backwards; since i cant remmeber how to do these..

OpenStudy (istim):

Well, it's reall just y=mx+b that we're looking at, with slope nad b sought for.

OpenStudy (anonymous):

f(x)=270+6(x/1.2)?

OpenStudy (anonymous):

but this isn't continuous function is it? you cant' sell 270.3333 cans can you?

OpenStudy (istim):

The answer in the textbook is 86-0.2x.

OpenStudy (anonymous):

the function?

OpenStudy (istim):

Yes.

OpenStudy (anonymous):

shouldn't the constant term be 270 not 86. as you start from the 270 figure not 86... the negative sign might have been because of the decrease.